HUBBLE SPACE TELESCOPE DAILY REPORT #2254 PERIOD COVERED: 0000Z (UTC) 11/18/98 - 0000Z (UTC) 11/19/98 Daily Status Report as of 323/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 8086 (Leonid STIS proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of quasar PKS2200-23 during the Leonid meteor shower (HST Leonid-avoidance pointing) to identify high redshift galaxies associated with Lyman-alpha and metal absorption systems in the quasar spectrum. Our key aim is to achieve the direct optical identification of high redshift galaxies which produce absorption against the background QSO. The observations were completed as scheduled, and no anomalies were noted. 1.2 Completed Two Sets of NIC/1/2/3 7961 (Flats: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the flat fields during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. These flats will provide a monitor for particulate comtamination and detector lateral position {from the coronagraphic spot and FDA vignetting}. The proposal completed with no reported problems. 1.3 Completed WF/PC-2 8087 (Leonid WF/PC-2 proposal) The WF/PC-2 was used to make observations during the Leonid meteor shower. This is a special pure-parallel service observation proposal which will execute during the HST Leonid-avoidance pointing. A deep WF/PC-2 image was taken over 6 orbits {in parallel with a STIS pointed observation}. Since HST was pointed toward the anti-Earth-vector, the image will be especially valuable in searching for slow-moving Kuiper belt objects. It will also be useful for high-redshift galaxy studies. The observations were completed as planned, and no anomalies were noted. 1.4 Completed Eight Sets of NIC/1/2/3 7963 (Darks: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. The observations were completed as planned, and no anomalies were noted. 1.5 Completed STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed Two Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.7 Completed NIC/1/2/3 8088 (Cosmological Parameters Omega and Lambda from High-Z Type Ia Supernovae - DD) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of a high-Z supernova in order to determine the cosmological parameters Omega and Lamda. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed WF/PC-2 7466 (From Globular Clusters to Tidal Dwarfs: Structure Formation in Tidal Tails.) The WF/PC-2 was used to make observations of the globular clusters (et al.) within the tidal tail of galactic merger NGC-7252. By modeling these systems with N--body hydrodynamic codes, can we identify the physical circumstances and formation histories which can produce tidal dwarfs, star clusters, and other structures in tidal tails. The observations were completed as planned, and no anomalies were reported. 1.9 Completed NIC/1/2/3 6482 (Weak Lensing in the Field of Luminous Quasars) The Near Infrared Camera (Cameras One, Two, and Three) was used to observe the relatively bright QSO PKS0135-24 which has long been suspected to be gravitationally magnified. We have found along the line of sight towards some of them a coherent shear pattern of faint distant galaxies. Moreover there appear to be overdensities of galaxies concentric with the shear patterns which are probably responsible for the weak lensing effect and the QSO magnifications. The observations completed as planned, with no reported anomalies. 1.10 Completed Five Sets of STIS/CCD 8062 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations taken in parallel with the prime science observations. The observations were completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 11 Successful : 11 2.2 FHST Updates: Scheduled : 10 Successful : 10 2.3 Operations Notes: The SSR EDAC error counter was cleared twice per ROP SR-1. The PSEA was re-configured to not close the aperture door (upon entry into hardware sunpoint) at 322/1728Z, the pre-Leonid Shower state. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH